# Testing Hypotheses of Covariate Effects on Topics of Discourse

### v1

@@ 3.2 Beta Regression
    Again focusing on attention on $\bm{\Theta}$ (as extracted by algorithm (1)), recall that the entries $\theta_{ij}\in[0,1]$ describes the probability (or “weight”) assigned to topic $i$ in document $j$. After selecting topic $i$ of interest to the analyst, and normalizing $\widetilde{\theta}_{ij}=\theta_{ij}/\sum_{i}\theta_{ij}$ so that rows sum to 1, vector $\widetilde{\theta}_{i}$ contains elements interpretable as the probability a word from document $j$ as being from topic $i$. For a fixed $i$, modelling $\theta_{ij}$ as depending on a vector of covariates $\mathbf{z}^{j}$ through coefficient vector $\bm{\beta}_{i}\in\mathbb{R}^{P}$,
    $\widetilde{\theta}_{ij}\sim Beta(\mbox{mean}=\mu,\mbox{precision}=\sigma),with$
    $g_{\mu}(\mu)=\mathbf{B}_{\mu}^{\text{T}}\mathbf{z}^{j},\mbox{ and }g_{\sigma}(\sigma)=\mathbf{B}_{\sigma}^{\text{T}}\mathbf{z}^{j}.$
    The mean-precision parameterization of the Beta provides a more intuitive interpretation of regression (Vasconcellos and Cribari-Neto, 2005). The link functions, $g_{\mu}(\mu)$ and $g_{\sigma}(\sigma)$, provide mappings from potentially different subsets of $\mathbf{z}^{j}$ into the appropriate spaces with covariate driven variation (Simas et al., 2010).
@@ 4 Experiments and Applications
    This section presents case studies showcasing BRETT. The first analysis considers the time evolution of discourse in papers from the NeurIPS conference to compare STM and BRETT. BRETT is shown to be much faster than the STM framework while retaining inferential accuracy. The second analysis considers the regional differences in vocabulary used to describe beer flavours.
@@ 4.1 Comparisons with STMs
    The STM class of probabilistic models is that which is closest in spirit to BRETT. However, that STMs permit covariates to influence both topic content and topic prevalence. In BRETT-style terminology, both $\bm{\Theta}$ and $\bm{\Phi}$ are thought to depend on the specified design matrix. One could compute regression-style statistics using $\bm{\Phi}$ (akin to what was accomplished in the previous section) to mimic this ability, but the efficient bootstrap-based error bars would be lost, as one would need to re-compute $\bm{\Phi}$ with each bootstrap sample. Accordingly, we only compare to those aspects of STMs most directly comparable to BRETT: their topic discovery and incorporation of “$\bm{\Theta}$-influencing” covariates.
@@ 4.1.1 Topic Extraction
    Fixing the number of topics $T=100$, tables (6) and (7) show each method’s ability to extract topics from the data. We make no claim as to 100 being the “correct” number of topics, but this choice appears to give sensible results in practice. Similarly, we elect not to compute any metrics measuring the quality of these topics. Rather, we are content to say that both methods produce what look to be very semantically coherent topics and are thus defensible methods for vanilla topic modelling tasks. That said, BRETT does have several interpretability advantages. First, anchor words provide “labels” of each topic, though this point seems moot in this example: the topics found by both methods appear easily understandable to humans. More importantly, BRETT ranks topics using the $\lambda$-criterion described in section (LABEL:sec:nmf), a feature lacking in PTMs that emphasizes how predictive the anchor is of its constituent words. BRETT ranks specific, semantically interesting topics higher.
    Although the topics differ outside of their leading words, this ensures that they are at least somewhat semantically comparable; in the case of the “Gaussian” topic, tables (6) and (7) in appendix A suggest that both methods are referring primarily to Gaussian processes.
    Data fits are shown in Figure 1. The STM results are from OLS and Beta regression as estimated by the mean of 10,000 draws from the variational posterior and computing the associated regression coefficients each time. The BRETT results are from fitting Beta regression directly and from the mean of 10,000 bootstrap OLS samples. Results are similar owing to the discrete nature of the covariates.
    Uncertaintly characterization differs in interpretation across the methods. STM tools and BRETT OLS sample the regression surface producing uncertainty intervals for the model fit. BRETT Beta instead fits a Beta distribution to the data, naturally producing prediction intervals for new observations. This difference characterizes the differing widths of intervals overlayed in Figure 1.
@@ 4.1.2 Computational Speed
    Using the NeurIPS data set, STMs and BRETT were compared for computational speed using the following experimental setup. The number of topics $T$ was chosen to range over the values $\{10,50,100,200\}$ and a binary variable $C\in\{T,F\}$ was devised, representing whether covariates should be included in the analysis or not. If not, classical covariate-free topic modelling was performed. Otherwise, covariate effects were estimated and 1000 simulations computed to assess the uncertainty in these estimates. In the case of BRETT, this means 1000 bootstrap samples were drawn. For the STM, “simulations” refers to 1000 draws from the posterior, which are then passed to the package’s estimateEffect(...) function. This function fits a linear model between the topic proportions (as drawn from the posterior) and covariates. STMs can be fit using several initialization schemes; we use the recommended default. Interestingly, this fits an anchor word-based PTM as a starting value for the variational inference scheme. All $4\times 2=10$ unique settings of $T$ and $C$ were timed using the same software on the same machine; figure (2) shows the results. BRETT is typically an order of magnitude faster and incurs minimal slowdowns when incorporating the extra regression functionality.
@@ 4.1.3 Inferential Accuracy
    The simulation study is performed by drawing 1000 TDMs by sampling words based on the word within topic probabilities and the topic allocations based on the continuous covariate. To reduce lack of identifiability, the anchors are fixed to those used to simulate the TDMs. Given the simulation mechanism, this study explores the variability with respect to changes in $N_{D}$, and $D$ with fixed anchors, using both recalculated $\Phi$ specific to each simulation or fixing $\Phi$ at the value used to simulate the TMDs.
    To ensure a consistent simulation setup, the simulation details are somewhat atypical. First the TDM is sparsely populated, moving through the list of words and picking a random document for each in which to place a single word count. This ensures the TDMs retain the same dimension at each iteration by forcing every word to appear in at least one document under all simulations. Ensuring that all words appear at least once in the corpus alters the original probabilities of both topics within documents and terms within topics. Consequently, regression coefficients estimated from this simulation study mechanism will not be directly comparable to the ground truth used to construct the TDM. To assess the stability of the regression coefficient estimates, a pseudo-ground truth must be devised. A pseudo ground truth is constructed by summing the 1000 TDMs elementwise into the equivalent of a TMD where each document has $1000\times N_{D}$ words per document.
    The alternatives were to elimiate low probability words resulting in a much smaller vocabulary or allow a varying vocabulary size across simulations. These alternative simulation designs are also expected to alter the covariate effects from their simulated targets. The Mean Squared Error for the covariate effect effect is shown in figure 3.
    Figure 3: Mean Squared Error in estimating regression effects using Beta regression with respect to the number of documents (rows) and number of words per document while recalculating $\Phi$ for each sample (left) or using a single fixed $\Phi$ for all simulations (right). Mean Squared Error in estimating regression effects using Beta regression with respect to the number of documents (rows) and number of words per document while recalculating $\Phi$ for each sample.
@@ 5 Application: The “Beer Data”
    A dataset about Canadian beer reviews is detailed in section 5.1 is used to show the method is able to extract well known results about flavours that define different beer styles in section 5.2. Section 5.3 tests hypotheses about whether local production of key ingredients has an affect on discourse around beer flavours. The section concludes with an examination of the stability of results under different scenarios.
@@ 5.2 Beer Style Hypotheses
    The dataset was filtered down to 7 beer styles, counts of which are given in table 1. The NMF model was run with 50 topics.
    The ability to test hypotheses depends on having a good metric for the characteristic of interest. The 50 topics, were examined manually to find measures of flavour attributes that coincide with the characteristics of the beer style definitions (Palmer, 2017). Some flavour attributes were spread into multiple topics, similar to the way lab sciences often indirectly measure attributes of interest from different perspectives and experiments.
@@
    The main terms composing the topics to test hypotheses are given in table 2. For the sake of brievity the table shows the most relevant terms from the top 20 most important terms per topic.
    The hypotheses are tested through logistic link function in Beta regression. Regression coefficients for hypotheses (H1-H8) are shown in table 3 as the deviation from the baseline beer style. In all cases our choice of baseline has a strong topic effect. Deviations from the baseline are substantial and negative except in ways that coincide with expected definitions and closely related beer styles. Imperial IPAs are considerably stronger than IPAs and often darker and with more hops. However the flavour profiles have considerable overlap compared to the within style variability in topics tropical (H1 p-value 0.08), grapefruit (H2 p-value 0.34)), and pine (H3 p-value 0.94). Porters and stouts are both made from dark roasted and chocolate malts resulting in substantial overlap in flavour compared to variability within styles leading to no significant difference in roasted (H4 p-value 0.25) or chocolate (H6 p-value 0.88) topics. Coffee complements the rich malty flavours of dark ales, but is more commonly added to stouts than in porters, resulting in a small but significant difference between beer styles in coffee (H5, p-value 0.0037). The dark malt base provides a creaminess, often enhanced by oatmeal (in stouts in particular), or the addition of nitrogen carbonation. Pumpkin ales also have creaminess from their pie-like flavours appearing as similarity to baselines (H7 difference between stouts and porters p-value .26, and difference between stouts and pumpkin ales p-value 0.070). Saisons are considerably yeastier in flavour than other beers, the closest being lagers and pumpking ales (H8 difference between saisons and lagers p-value 2e-36, and difference between saisons and pumpkin ales p-value 8e-10). Lagers are generally light in flavour and sometimes the yeast is a dominant flavour, but far from the prominance of a saison. Pumpkin ales have considerable variability spanning and could be modified from anything from a saison to a stout base. These effects are all reflected in the prediction intervals of figure 4 for topics in table 2. Beta Distributions have full support over $[0,1]$ leaving effect size prediction intervals wide with some strong prediction confidence close to 0, even when an overall effect is large as is seen in all the baseline beer styles for the hypotheses.
@@
    grapefruit pineapple, fresh, crisp, love, citrusy, backbone, excellent, delicious, juicy, lacing, resiny, clean
    tropical mango, pineapple, fruits, lemon, gold, tangerine, dank, juicy, rind, citra, peach, papaya
    pine piney, resin, fresh, hoppy, balance, oily, west coast, balanced, huge, sticky, plenty, dipa
    yeast spice, white, pepper, lemon, spicy, wheat, golden, belgian, yellow, yeasty, banana, fruity
    brett funk, tart, funky, oak, sour, wine, tartness, lemon, apple sourness, barnyard, wild
    pumpkin cinnamon, spices, nutmeg, spice, pie, ginger, clove, amber, allspice, copper, fall, flavour
    roasted black, burnt, roasty, oatmeal, tan, thick, smooth, oats, sweetness, opaque, brown, espresso
    coffee espresso, strong, beans, brown, mouthfeel, opaque, drinkability, bodied, heavy, hints, retention, barley
    chocolate cocoa, milk, mocha, hint, black, smooth, mouthfeel, roast, characters, delicious, lacing, chocolatey
    creamy smooth, thick, brew, excellent, rich, cream, bodied, dense, sweetness, nitro, silky, feel
@@
    H1 - tropical 0.28 NS -2.77$\star$ -3.07 -4.19 -4.90 -1.39 -3.75
    H2 - grapefruit -0.15 NS -2.75$\star$ -2.57 -4.94 -4.27 -2.20 -8.39
    H3 - pine 0.01 NS -3.35$\star$ -2.56 -4.52 -5.48 -2.06 -4.41
    H4 - roasted -5.10 -4.10 -3.95 -0.18NS -2.36 -3.46 -2.08$\star$
    H5 - coffee -6.17 -4.53 -5.05 -0.42 -3.83 -4.93 -2.48$\star$
    H6 - chocolate -6.32 -4.84 -6.77 -0.02NS -3.03 -4.89 -2.83$\star$
    H7 - creamy -1.05 -0.92 -1.59 -0.22NS -0.62NS -1.38 -3.36$\star$
    H8 - yeast -2.21 -2.00 -1.95 -2.68 -2.26 -1.35$\star$ -2.88
@@ 5.3 Regional Differences in Flavour Characteristics of IPAs
    Interest lies in assessing if there are effects of ingredient production within the flavour category. Specifically we are interested in knowing if the Prairies have grainier, breadier IPAs or if BC and Ontario have hoppier IPAs owing to localized production of base ingredients. Hypothesis tests for regional effects on flavour discourse will be confounded with differences in the types and numbers of beers produced. Focusing within a single product category, IPAs are used to test hypotheses that there are regional differences in flavour discourse. The NMF model is refitted using 776 IPAs. Using too general a corpus of all beer styles to devise topics results in messier topics that are less interpretable or useful (see for example Tshitoyan et al. (2019), Luca et al. (2022)). NMF was fitted with 50 topics resulting in topics that interpretably showcase the flavours of interest for these hypotheses.
@@
    The grainy, bready flavour set can be assessed using earthy (H9) and caramel (H10) topics while the hoppier characteristics were assessed using grapefruit (H11) and tropical (H12). Table 4 provides characteristic terms within these topics.
    The probability of topic $Y_{i}$ for the $i^{th}$ beer is modelled with an intercept effect for the region of interest with other covariates representing the $j^{th}$ regional deviation from the intercept.
@@
    Regression coefficients for hypotheses (H9-H12) are shown in table 4. The Prairies show significantly more earthy and caramel flavour discourse than any other region. These topics represent flavours that are not typical of the IPA style, suggesting an influence of locally produced ingredients. The grapefruit and tropical topics describe different dimensions of hop flavours characteristic of IPAs. The vast majority of beers are brewed with dried, vacuum sealed hops that travel well with reasonable shelf life. It should be expected that discourse around all IPAs mention these topics with considerable probability. It is not surprising that there are almost no significant regional differences from the BC baseline. The exception is that Ontario is showing a somewhat positive difference from BC (H11 difference between BC and Ontario p-value 0.019) in discourse of the grapefruit topic characterized by crisp, citrusy, resiny terms. Prediction intervals for effects are shown in figure 5.
    5.4 Stability With Respect to Number of Topics
    When the procedure in Section 5.2 was performed using 150 topics, the results were largely the same since additional anchors extend the topics into a higher dimensional orthogonal space. The utility of new topics starts to break down when too many topics are added because each new topic chips away at existing topics. Eventually a topic may split into two, diffusing the covariate of interest and rendering the hypothesis more challenging to define. Note that the probability of a topic for a document is monotone decreasing in the number of topics meaning that results are conditional on the model structure.
    5.5 Stability With Respect to General vs Curated Corpus
    When the topics in Section 5.3 were formed using the entire beer dataset, the topics were tuned towards separating out the broad differences, mostly driven by different beer styles. When using those topics and subsetting the documents to just IPAs, the flavour effects of interest were too diffuse to be usefully measured. Language models have been shows to have the best success when trained on data tuned to the task of interest (Tshitoyan et al., 2019).
@@ 5.5 Stability With Respect to General vs Curated Corpus
    Table 4: Topics used to test flavour attribute hypotheses H9-H12 by region within the IPA beer style.
@@
    Figure 5: Prediction interval for flavour topic probability for different regions using 50 topics.
    6 Conclusion

### v2

@@ 3.2 Beta Regression
    Again focusing on attention on $\bm{\Theta}$, as extracted by algorithm 1, recall that the entries $\theta_{ij}\in[0,1]$ describe the probability (or “weight”) assigned to topic $i$ in document $j$. After selecting topic $i$ of interest to the analyst, and normalizing $\widetilde{\theta}_{ij}=\theta_{ij}/\sum_{i}\theta_{ij}$ so that rows sum to 1, vector $\widetilde{\theta}_{i}$ contains elements interpretable as the probability of a word from document $j$ as being from topic $i$. Again, the decomposition $\bm{\Theta}=\left(\bm{\Lambda}\right)^{-1}\mathbf{X}_{\dagger}$, followed by the rescaling in $\widetilde{\theta}$ highlights that Beta regression is conditional only on the anchor block. For a fixed $i$, modelling $\widetilde{\theta}_{ij}$ as depending on a vector of covariates $\mathbf{z}^{j}$ through coefficient vector $\bm{\beta}_{i}\in\mathbb{R}^{P}$,
    $\widetilde{\theta}_{ij}\sim Beta(\mbox{mean}=\mu,\mbox{precision}=\sigma).$ (6)
    The mean-precision parameterization of the Beta regression model provides a more intuitive interpretation of regression (Vasconcellos and Cribari-Neto, 2005). The link functions map potentially different subsets of $\mathbf{z}^{j}$ into the appropriate spaces with covariate driven variation (Simas et al., 2010)
    $g_{\mu}(\mu)=\mathbf{B}_{\mu}^{\text{T}}\mathbf{z}^{j},\mbox{ and }g_{\sigma}(\sigma)=\mathbf{B}_{\sigma}^{\text{T}}\mathbf{z}^{j}.$ (7)
@@ 4 Experiments and Applications
    This section presents case studies showcasing BRETT. The first analysis considers the time evolution of discourse in papers from the NeurIPS conference to compare STM and BRETT. A simulation example follows showcasing behaviour with respect to changes in document length and number.
@@ 4.1 Comparisons with STMs
    The STM class of probabilistic models is that which is closest in spirit to BRETT. However, that STMs permit covariates to influence both topic content and topic prevalence. In BRETT-style terminology, both $\bm{\Theta}$ and $\bm{\Phi}$ are thought to depend on the specified design matrix. One could compute regression-style statistics using $\bm{\Phi}$ (akin to what was accomplished in the previous section) to mimic this ability, but the efficient bootstrap-based error bars would be lost, as one would need to re-compute $\bm{\Phi}$ with each bootstrap sample. Accordingly, we only compare to those aspects of STMs most directly comparable to BRETT: their topic discovery and incorporation of $\bm{\Theta}$-influencing” covariates.
@@ 4.1.1 Topic Extraction
    Fixing the number of topics $T=100$, tables (7) and (8) show each method’s ability to extract topics from the data. We make no claim as to 100 being the “correct” number of topics, but this choice appears to give sensible results in practice. Similarly, we elect not to compute any metrics measuring the quality of these topics. Rather, we are content to say that both methods produce what look to be very semantically coherent topics and are thus defensible methods for vanilla topic modelling tasks. That said, BRETT does have several interpretability advantages. First, anchor words provide “labels” of each topic, though this point seems moot in this example: the topics found by both methods appear easily understandable to humans. More importantly, BRETT ranks topics using the $\lambda$-criterion described in section 2.2, a feature lacking in PTMs that emphasizes how predictive the anchor is of its constituent words. BRETT ranks specific, semantically interesting topics higher.
    Although the topics differ outside of their leading words, this ensures that they are at least somewhat semantically comparable; in the case of the “Gaussian” topic, tables 7 and 8 in appendix A suggest that both methods are referring primarily to Gaussian processes.
    Regression fits to the data for the are shown in Figure 1 using binned time as the discrete regression covariates. The STM results are from OLS and Beta regression as estimated by the mean of 10,000 draws from the variational posterior and computing the associated regression coefficients each time. The BRETT results are shown fitting Beta regression directly and also from the mean of 10,000 bootstrap OLS samples. Results are similar owing to the discrete nature of the covariates.
    Uncertainty characterization differs in interpretation across the methods. STM and BRETT OLS sample the regression surface producing uncertainty intervals for the model fit. BRETT Beta regression instead fits a Beta distribution to the data, producing prediction intervals for new observations. This difference in interpretation is highlighted in the differing widths of intervals overlayed in Figure 1.
@@ 4.1.2 Computational Speed
    Using the NeurIPS data set, STMs and BRETT were compared for computational speed using the following experimental setup. The number of topics $T$ was chosen to range over the values $\{10,50,100,200\}$ and a binary variable $C\in\{T,F\}$ was devised, representing whether covariates should be included in the analysis or not. If not, classical covariate-free topic modelling was performed where STM and BRETT reduce to LDA with NMF respectively. Otherwise, covariate effects were estimated and 1000 simulations computed to assess the uncertainty in these estimates. In the case of BRETT, this means 1000 bootstrap samples were drawn. For the STM, “simulations” refers to 1000 draws from the posterior, which are then passed to the package’s estimateEffect(...) function. This function fits a linear model between the topic proportions (as drawn from the posterior) and covariates. STMs can be fit using several initialization schemes; we use the recommended default. Interestingly, this fits an anchor word-based PTM as a starting value for the variational inference scheme. All $4\times 2=10$ unique settings of $T$ and $C$ were timed using the same software on the same machine; figure (2) shows the results. BRETT and NMF are typically an order of magnitude faster and incurs minimal slowdowns when incorporating the extra regression functionality.
@@ 4.2 Inferential Accuracy
    The simulation study is performed by drawing 1000 TDMs by sampling words based on the word within topic probabilities and the topic allocations based on the continuous covariate. To reduce lack of identifiability, the anchors are fixed to those used to simulate the TDMs. Given the simulation mechanism, this study explores the variability with respect to changes in $N_{D}$, and $D$ with fixed anchors, using both recalculated $\Phi$ specific to each simulation or fixing $\Phi$ at the value used to simulate the TMDs. Recalculating $\Phi$ for each dataset is equivalent to having an analyst select the anchor words and let the data fine tune the interpretation of the topic. Holding $\Phi$ fixed at the value used to construct the simulation defines $\Theta=(\Phi^{\prime}\Phi)^{-1}\Phi^{\prime}X$ and is equivalent to the analyst pre-defining topics from an external independent dataset and uses those topic definitions to define document topic weights.
    To ensure a consistent simulation setup, the simulation details are somewhat atypical. First the TDM is sparsely populated, moving through the list of words and picking a random document for each in which to place a single word count. This ensures the TDMs retain the same dimension at each iteration by forcing every word to appear in at least one document under all simulations. Ensuring that all words appear at least once in the corpus alters the original probabilities of both topics within documents and terms within topics. Consequently, regression coefficients estimated from this simulation study mechanism will not be directly comparable to the ground truth used to construct the TDM. To assess the stability of the regression coefficient estimates, a pseudo-ground truth must be devised. A pseudo ground truth is constructed by summing the 1000 TDMs element-wise into the equivalent of a TMD where each document has $1000\times N_{D}$ words per document.
    The alternatives were to eliminate low probability words resulting in a much smaller vocabulary or allow a varying vocabulary size across simulations. These alternative simulation designs are also expected to alter the covariate effects from their simulated targets. The Mean Squared Error for the covariate effect effect is shown in figure 3. The strategy of recalculating $\Phi$ for each TDM while holding the anchors fixed performs well with the MSE decreasing quickly in both number of documents and number of words per document. Allowing $\Phi$ to be recalculated introduces some variability in the balance of words within topic, but overall seems to estimate a $\Phi$ that is close to its asymptotic baseline value. The strategy of recalculating $\Phi$ for each dataset results in around an order of magnitude decrease in MSE compared to using a fixed $\Phi$. In the fixed $\Phi$ strategy, $\Phi$ is held at the wrong value as it does not account for the simulation mechanism that ensures the vocabulary size remains fixed leading to a decrease in performance. This suggests that the analyst is best served by selecting anchor terms and letting the NMF routine define the allocation of terms into topics.
    Figure 3: Mean Squared Error in estimating regression effects using Beta regression with respect to the number of documents (rows) and number of words per document while recalculating $\Phi$ for each sample (left) or using a single fixed $\Phi$ for all simulations (right).
@@ 5 Application: The “Beer Data”
    Collection and processing of the Canadian beer review dataset is detailed in section 5.1 and used to show BRETT is able to extract well known results about flavours that define different beer styles in section 5.2. Section 5.3 tests hypotheses about whether local production of key ingredients has an affect on discourse around beer flavours. The section concludes with an examination of the stability of results under different scenarios.
@@ 5.2 Beer Style Hypotheses
    The dataset was filtered down to 7 beer styles, counts of which are given in table 1. NMF was run with 50 topics.
    The ability to test hypotheses depends on having a good metric for the characteristic of interest. The topics were examined manually to find measures of flavour attributes that coincide with the characteristics of the beer style definitions (Palmer, 2017). Some flavour attributes were spread into multiple topics, similar to the way lab sciences often indirectly measure attributes of interest from different perspectives and experiments.
@@
    The main terms composing the topics to test hypotheses are given in table 2. For the sake of brevity the table shows the most relevant terms per topic. The hypotheses are tested using Beta regression from equation (6) where the design matrix $Z\in\{0,1\}^{D\times 7}$ is defined with categorical covariates corresponding to the 7 beer styles through link functions
    $g_{\mu}(\mu)=logit(\mu)=\mathbf{B}_{\mu}^{\text{T}}\mathbf{z}^{j},\mbox{ and }g_{\sigma}(\sigma)=logit(\sigma)=\mathbf{B}_{\sigma}^{\text{T}}\mathbf{z}^{j}.$ (8)
    In each case one beer style is chosen as a baseline (intercept) and all other covariates are interpreted as deviations thereof for testing the null hypothesis of no difference in discourse about a flavour characteristic compared to the baseline beer style.
    Regression coefficients for hypotheses (H1-H8) are shown in table 3 as the deviation from the baseline beer style. In all cases our choice of baseline has a strong topic effect. Deviations from the baseline are substantial and negative except in ways that coincide with expected definitions and closely related beer styles suggesting a reduction in discourse associated with changing beer style. Imperial IPAs are considerably stronger than IPAs and often darker and with more hops. However the flavour profiles have considerable overlap compared to the within style variability in topics tropical (H1 p-value 0.08), grapefruit (H2 p-value 0.34)), and pine (H3 p-value 0.94). Porters and stouts are both made from dark roasted and chocolate malts resulting in substantial overlap in flavour compared to variability within styles leading to no significant difference in roasted (H4 p-value 0.25) or chocolate (H6 p-value 0.88) topics. Coffee complements the rich malty flavours of dark ales, but is more commonly added to stouts than in porters, resulting in a small but significant difference between beer styles in coffee (H5, p-value 0.0037). The dark malt base provides a creaminess, often enhanced by oatmeal (in stouts in particular), or the addition of nitrogen carbonation. Pumpkin ales also have creaminess from their pie-like flavours resulting in a similarity with baselines (H7 difference between stouts and porters p-value .26, and difference between stouts and pumpkin ales p-value 0.070). Saisons are considerably yeastier in flavour than other beers, the closest being lagers and pumpking ales (H8 difference between saisons and lagers p-value 2e-36, and difference between saisons and pumpkin ales p-value 8e-10). Lagers are generally light in flavour and sometimes the yeast is a dominant flavour, but far from the prominance of a saison. Pumpkin ales have considerable variability as they could be modified from nearly any beer style. These effects are all reflected in the fitted distribution of figure 4 for topics in table 2.
@@
    tropical mango, pineapple, fruits, lemon, gold, tangerine, dank, juicy, rind, citra, peach
    grapefruit pineapple, fresh, crisp, love, citrusy, backbone, excellent, delicious, juicy, lacing, resiny
    pine piney, resin, fresh, hoppy, balance, oily, balanced, huge, sticky, plenty
    roasted black, burnt, roasty, oatmeal, thick, smooth, oats, sweetness, opaque, brown, espresso
    coffee espresso, strong, beans, brown, mouthfeel, opaque, drinkability, bodied, heavy, hints, barley
    chocolate cocoa, milk, mocha, hint, black, smooth, mouthfeel, roast, characters, lacing, chocolatey
    creamy smooth, thick, brew, rich, cream, bodied, dense, sweetness, nitro, silky, feel
    yeast spice, white, pepper, lemon, spicy, wheat, golden, belgian, yellow, yeasty, banana
@@
[…]
    H2 - grapefruit -0.15 NS -$\star -2.57$ -4.94 -4.27 -2.20 -8.39
[…]
@@ 5.3 Regional Differences in Flavour Characteristics of IPAs
    Interest lies in assessing if there are geographic effects relating to proximity of ingredient production within their flavour categories. Specifically we are interested in knowing if the Prairies have grainier, maltier IPAs or if BC and Ontario have hoppier IPAs owing to localized production of base ingredients. Hypothesis tests for regional effects on flavour discourse will be confounded with differences in the types and numbers of beers produced, so we focus within a single product category, IPAs are used to test hypotheses. NMF is refitted using 776 IPAs (and Imperial IPAs) with 50 topics.
@@
    The grainy and malty flavours can be assessed using earthy (H9) and caramel (H10) topics while the hoppier characteristics are assessed using grapefruit (H11), tropical (H12), and pine (H13). Table 4 provides characteristic terms within these topics.
    Hypotheses are tested through logistic link function in Beta regression using equations (6) and (8) where the design matrix $Z\in\{0,1\}^{D\times 5}$ is defined with categorical covariates corresponding to the 5 regions. In each case one region style is chosen as a baseline (intercept) and all other covariates are interpreted as deviations thereof associated with changing regions. Under the null hypothesis of no difference between regions, we expect to observe a significant effect for the baseline while coefficients for all other regions take values of approximately 0.
@@
    Regression coefficients for hypotheses (H9-H13) are shown in table 5. The Prairies show significantly more earthy and caramel topic discourse than any other region. These topics represent flavours that are not typical of the IPA style, potentially influenced by local production of ingredients. The grapefruit, tropical, and pine topics describe different dimensions of hop flavours characteristic of IPAs. Almost no significant regional differences from the BC baseline were found. The exception is that Ontario is showing a somewhat positive difference from BC (H11 difference between BC and Ontario p-value 0.019) in discourse of the grapefruit topic characterized by crisp, citrusy, resiny terms. The vast majority of beers are brewed with dried, vacuum sealed hops that travel well with reasonable shelf life explaining the lack of overall flavour effects associated with proximity to hop farms. Predicted distribution of topic discourse for effects are shown in figure 5.
    Figure 5: Predicted distributions for flavour topics by region.
@@
    pine piney, resin, balance, huge, hoppy, oily, rind, balance, complex, lingering
    Table 4: Topics used to test flavour attribute hypotheses H9-H13 by region within the IPA beer style.
@@
    5.4 Stability With Respect to Number of Topics
    When the procedure in Section 5.2 was performed using 150 topics, the results were largely the same since additional anchors project the documents into a higher dimensional orthogonal space. The utility of new topics starts to break down when too many topics are added. Eventually a useful topic will split into two, diffusing the covariate of interest and rendering the hypothesis more challenging to define. Note that the probability of a topic for a document is monotone decreasing in the number of topics meaning that results are conditional on the model structure.
    5.5 Stability With Respect to General vs Curated Corpus
    When the topics in Section 5.3 were formed using the entire beer dataset rather than the IPA subset, the topics that emerged were tuned towards separating out the broad style differences since these captured most of the variation in the dataset. The resulting flavour topics were overly general when a small number of topics was used and when larger numbers of topics were used, topics of interest fractured off of the general topics in such a way that necessitated a larger number of hypotheses to test for the flavours of interest. Language models have similarly been shown to have the best success when trained on data tuned to the task of interest (Tshitoyan et al., 2019; Luca et al., 2022).
    5.6 Stability with Respect to Anchor Selection Variability
    The SPA routine chooses anchors that are highly used and orthogonal to those already selected. The analyst looks through the resulting anchors and the topics they define to assess their utility in measuring the pre-specified, hypothesized relationships. As a result the anchors selected by SPA (and used for our hypotheses) may vary if a new text sample is obtained. However this is different from the usual selection uncertainty problem which is cast as follows. Consider selecting an active subset of $x\subset X$ in regression $Y=f(X\beta)+\epsilon$. Here one might use LASSO or a stepwise routine which then requires post selection inference to produce interval estimates for the corresponding $\beta$ elements that properly account for the model selection. Our problem is inherrently different in that we are selecting appropriate values of $y\subset Y$ that suit our externally generated hypotheses. When performed using SPA, this action is taken without consideration of $X$. In text modelling all pre-processing steps impact the analytic outcome, whether it is converting words to lowercase, fixing typos, handling n-grams or stopwords. NMF decomposition and selection of anchors is a form of pre-processing step to be assessed for quality by the analyst taken before modelling begins. Analysis is then performed conditional on these decisions. Pre-processing is performed prior to considering covariate effects.
    In the absence of a particular anchor word, an important topic will not vanish, but instead will evolve through a different anchor label and perhaps a different number of meaningful topics for the hypothesized relationships. In an extreme demonstration of the impact of changing anchors, we reconsider the problem of assessing a regional effect on hoppiness (H11 - H13) from Section 5.3, while forcing SPA to exclude {grapefruit, tropical, pine} from being considered as anchors. As before, the analyst must examine the resulting topics and assess their validity towards the inferential goal. In this experiment, we find the replacement hoppiness topics to be less focused. The important terms within selected anchors {citrus, fruit, orange, piney} are listed in table reftable:50topicexclusions. As before we find no significant deviation from the British Columbia baseline at the 5% level. A more thorough exploration of any potential impact of pre-processing is left for future work.
    topic defining terms
    citrus pine hoppy smooth rind appearance grapefruit peel
    fruit tropical juicy mango apricot peach passion guava
    orange grapefruit pine citrusy lingering peel crisp oily
    piney pine grapefruit hoppy resin oily smooth copper
    Table 6: Topics used to re-test regional differences in hoppyness discourse while excluding the selection of anchors grapefruit, tropical, pine.
    6 Discussion